OpenStudy (anonymous):

A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound? A. 15 lbs oatmeal, 15 lbs bran flakes B. 20 lbs oatmeal, 10 lbs bran flakes C. 15 lbs bran flakes, 15 lbs oatmeal D. 20 lbs bran flakes, 10 lbs oatmeal

OpenStudy (mertsj):

Let x = amount of oatmeal 30-x= amount of bran flakes 3.6x+4.8(30-x)=4(30)

OpenStudy (anonymous):

Thank you :D

OpenStudy (mertsj):

yw